OpenStudy (anonymous):

Rationalize the denominator of √-36 --------- (2-3i)+(2+3i)

OpenStudy (anonymous):

well the denominator is imaginary (or complex) in its current state, but to make it rational, all you do is multiply it out. The root of -36 on the numerator is comlex too, but you want to rationalise the denominator so this need not be touched...

OpenStudy (anonymous):

oh it's a plus! my bad

OpenStudy (anonymous):

denominator is 4

OpenStudy (anonymous):

\[\frac{i\sqrt{6}}{4}\] does the job since here \[4\in \mathbb{Q}\]

OpenStudy (anonymous):

god damn i am stupid it is this \[\frac{3i}{2}\]

OpenStudy (anonymous):

the point is, 2 ∈ Q